OpenStudy (anonymous):

Find the length of a median of an equilateral triangle having sides of length 2. a) square root of 3 b) square root of 3/2 c) square root of 3/4 d) 3

Parth (parthkohli):

The median bisects(divides the side into two). Now, the median will make a 90 degree angle and will form two 30-60-90 triangles.  So, 2 will be 2x and the hypotenuse. 1 = x So, median would be \(1 \sqrt{3}\) = \(\sqrt{3}\)

Parth (parthkohli):

a) is the answer.

OpenStudy (shubhamsrg):

a median in an equilateral triangle actually intersects the oppotite side at right angles..so you have a right triangle with sides 2,1,and x..you can easily solve for x now..

OpenStudy (callisto):

median = height angle in a equi. triangle =60 use tan. tan60 = x / (2/2) x = tan60 =?

OpenStudy (shubhamsrg):

opposite*

Parth (parthkohli):

The ratio of a 30-60-90 triangle is 2x(hypotenuse), x(base), xsqrt3(perpendicular)

Parth (parthkohli):

I never answered better lol ;)